OpenStudy (anonymous):

Solve the equations: x+y=5, 2/x+3/y=2. . . . ANSWER = {(5/2,5/2),(2,3)}

OpenStudy (anonymous):

\[x+y=5 and \frac{ 2 }{ x }+\frac{ 3 }{y }=2\]

OpenStudy (anonymous):

answer\[[(\frac{ 5 }{2 },\frac{ 5 }{ 2 }),(2,3)]\]

OpenStudy (tyteen4a03):

The answers are correct. What are you trying to ask?

OpenStudy (anonymous):

how to solve this

OpenStudy (anonymous):

2/x +3/y = 2 Take L.C.M and you get, (2y+3x)/xy = 2 2y+3x = 2xy 2y+2x+x = 2xy 2(x+y)+x = 2xy 2(5)+x = 2xy 10+x = 2xy 10 = 2xy-x 10 = x(2y-1) -> (1) Now, x+y = 5 x = 5-y Substitute this expression in (1), you get 10 = (5-y)(2y-1) 10 = 11y-2y^2-5 Rearranging the terms, we get 2y^2-11y+15 = 0

OpenStudy (anonymous):

Now its a quadratic equation..Can you solve it for roots?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i did'nt understand this step 10 = 2xy-x 10 = x(2y-1) -> (1)

OpenStudy (anonymous):

I took out x as the common term from the right side terms...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Are you sure that you got it?

OpenStudy (anonymous):

yeah i am solving quadratic right now

OpenStudy (anonymous):

ok got it the answer is {(5/2,5/2),(2,3)}

OpenStudy (anonymous):

i solved it further after quadratic putting the values of y in x=5-y